OpenStudy (anonymous):

i need to put 9x-3y=-12 in slope intercept form

OpenStudy (anonymous):

can anyone help me?

OpenStudy (anonymous):

When you convert to slope intercept form, you're looking for y = mx+b in your equation. you have 9x, -3y, and -12 for b, solve for y and you'll get your answer.

OpenStudy (anonymous):

wats the m part?

OpenStudy (anonymous):

m is the slope, currently, you don't know the slope because of the -3y, when you solve for y, you will have to divide by three eventually, thus changing the current constant on x.

OpenStudy (anonymous):

how do i set it up to solve for y?

OpenStudy (anonymous):

You have to isolate y to have it alone on one side of the equals sign. -3y=-9x-12, this is the first step. Notice how 9x has now become negative, that is because we have subtracted 9x from both sides, to isolate y. Now to complete it, y must be positive, and y alone with no numbers on it. Can you figure it out?

OpenStudy (anonymous):

idk hold on lemme try

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

so do i add -12 to both sides?

OpenStudy (anonymous):

You don't have to, because y is already alone, when we're in slope intercept form, always remember the pattern y = mx+b where m is the slope, and b is the y intercept, hence the name.

OpenStudy (anonymous):

ohk so i have 3y=-9x+-12 right?

OpenStudy (anonymous):

then now im trying to get rid of the three.... but how? i divide?

OpenStudy (anonymous):

When you do something to one side of the equation, you must do it to the other, removing the negative from the y is as if you multiplied it by -1, which means you must multiply the other side by -1 as well, leaving 3y=9x+12 and yes, you divide by 3 in the end.

OpenStudy (anonymous):

ohk so my brain is not comprehending this im so sorry i hope im not aggrivating u... but ohk where do i go from 3y=9x+12?? im not understanding and i thank u for being patient nd helping me but is the answer y=3x=4? if it is how? because someone juss gave me the answer

OpenStudy (anonymous):

That's the correct answer if you meant + and not = between 3x and 4, basically. 9x - 3y =-12 subtract 9x from both sides -3y = -9x -12 multiply both sides by -1 3y = 9x +12 divide by 3 y = 3x +4 final answer You may even combine steps 2 and 3 by just dividing by -3, much faster, I just wanted to show it like that so you could see why it happens.

OpenStudy (anonymous):

omgee i get it now!! and yes thats wat i meant lol thanks so much for all ur help!!